# On the Within-class Variation Issue in Alzheimer’s Disease Detection

### v2

@@ Abstract
Alzheimer’s Disease (AD) detection employs machine learning classification models to distinguish between individuals with AD and those without. Different from conventional classification tasks, we identify within-class variation as a critical challenge in AD detection: individuals with AD exhibit a spectrum of cognitive impairments. Therefore, simplistic binary AD classification may overlook two crucial aspects: within-class heterogeneity and instance-level imbalance. In this work, we found using a sample score estimator can generate sample-specific soft scores aligning with cognitive scores. We subsequently propose two simple yet effective methods: Soft Target Distillation (SoTD) and Instance-level Re-balancing (InRe), targeting two problems respectively. Based on the ADReSS and CU-MARVEL corpora, we demonstrated and analyzed the advantages of the proposed approaches in detection performance. These findings provide insights for developing robust and reliable AD detection models.
@@
Alzheimer’s disease, neurocognitive disorder, within-class variations, AD detection, dementia, healthcare1 Introduction
Neurocognitive disorders (NCD) such as Alzheimer’s Disease (AD), present a substantial and growing challenge within the aging population, characterized by progressive cognitive decline across multiple domains, including memory, attention, and executive function [1]. For timely intervention and management, in-person clinical assessments have been the primary protocol for screening AD patients in populations, where participants are examined using specially designed assessment tasks to test potential abnormal declines in cognitive abilities[2, 3, 4]. In contrast to traditional on-set assessments, recent advancements in machine learning technologies have facilitated speech-based automatic AD detection as a promising screening approach, with the advantages of being scalable, accessible, and cost-efficient.
@@
Beyond the challenge of feature extraction and representation learning, we propose that AD detection faces unique difficulties inherent to the nature of Alzheimer’s Disease. Medical literature [21, 22, 23] establishes AD as a degenerative disorder characterized by a continuum of pathophysiological changes, resulting in gradual cognitive and functional decline. This spectrum of cognitive performance among AD patients is illustrated in Fig. 1(a). Consequently, under standard classification modeling, samples with the same label may exhibit varying degrees of cognitive impairment patterns, which could lead to considerable variability in recognition features. This distinguishes AD detection from classification tasks such as image classification, which typically involve more consistent features under clearly defined categories. An ideal solution would be to model AD detection by regression or multi-way classification to capture this continuum. However, granular labels are rarely provided for most cognitive assessment tasks, especially for open-source datasets. While reference continuous scores are occasionally available (such as Mini-Mental State Examination scores), these scores are obtained from different tests that are distinct from the assessment in the data and can hardly be regarded as gold standard for supervision. Therefore, binary classification remains the prevailing method for Alzheimer’s Disease (AD) detection.
This work takes a first step towards the within-class variation (WCV) issue in AD detection. We inspect the sample variability beyond binary labels using the English ADReSS[24] and Cantonese CU-MARVEL [6] datasets (see Fig. 1 (b) and Fig. 1 (c)). Even with comparable class members, AD samples functioning a larger variance in cognitive function compared to healthy samples. Given these findings, we propose that the conventional binary classification paradigm overlooks two critical aspects: a) within-class heterogeneity (WCH): samples with varying AD severity are assigned to the same class, potentially inhibiting the model’s sensitivity to certain changes in cognitive function; and b) instance-level imbalance (ILI): the frequency of varying-severity samples is imbalanced even with balanced class size, thereby introducing potential bias. A key challenge in tackling these issues is the lack of informative instance-level labels indicating sample severity. Accordingly, this work explored sample score estimation that distills proxy soft labels from ensemble models with only hard label supervision. Building upon sample score estimation, we subsequently propose two approaches addressing WCH and ILI respectively: Soft Target Distillation (SoTD) and Instance-level Re-balancing (InRe). SoTD approach leverages the label refinery approach[25] to train classifiers using informative soft label supervision. And InRe approach re-weights imbalanced samples at instance level using log posterior ratios and inverse kernel density.
@@ 4.1 Validate estimated sample scores
An effective sample score estimation serves as a prerequisite for modeling within-class heterogeneity and imbalance using SoTD and InRe approaches. We first generate estimated scores for every sample in two datasets and observe their correlation with participants’ cognitive scores (i.e., MMSE scores in ADReSS; MocA scores in CU-MARVEL). It is important to note that the cognitive scores referenced here should be viewed as assistant labels that indicate participants’ cognitive ability in certain aspects, whereas the hard labels represent ground-truth labels for AD. As drawn in Fig. 3, we can observe that although the cognitive score labels were not seen during training, the estimated sample scores exhibit alignment with cognitive scores. Their Pearson Correlation is and in the two datasets respectively. Correspondingly, as used in the proposed SoTD method, guiding classifier training with these soft targets would take into account more comprehensive cognitive abilities of participants compared to relying on binary hard labels. Although we found some soft targets show “noisy” deviations from cognitive scores, we contend that the model overall benefits from these soft targets rather than just binary labels with values of 1 and 0.
@@ 4.2 AD detection performance
InRe* | .5518 | .5881 |
Table 1 presents our experimental results on ADReSS and CUMV-R corpus. To reduce randomness, we performed 10-fold cross-validation for 20 random runs and reported the averaged results. On the ADReSS corpus, the SoTD approach outperformed the baseline and ensemble systems across all 4 metrics. This demonstrates general improvement in AD classification. The InRE approach also led to improvement in the balanced accuracy and F1 metrics, while remarkably enhancing the recall rates. This aligns with our expectation because the InRe approach could emphasize the infrequent positive (AD) samples in the dataset (as shown in Fig. 1 (b)), thereby improving the model’s sensitivity. It is worth noting that this contribution differs from conventional class re-balancing methods since the number of positive and negative samples is equal in the training set. Thus, we contend that the advantage of the InRe approach stems its awareness of within-class imbalance. The resampling approach was not carried out in this dataset, as this dataset is already balanced and the results will be the same as the baseline.
The CUMV-R dataset poses a more challenging scenario for several reasons: (a) it contains more variations in the cognitive scores; (b) positive and negative samples are more imbalanced (ratio=1.51); and (c) Cantonese BERT models might not be as well-developed as English models. This is particularly evident in the low recall rates. The SoTD approach outperform both baseline and ensembling methods by balancing the model’s predictions, evidenced by the improved recall rate. The resampling and InRe methods both show further improvement in recall rate, while the InRe method improved the models’ sensitivity by a large margin () and therefore remarkably improved the overall F1 score ().
@@ 4.3 Inspect sample weights in InRe approach
To gain deeper insight into the InRe approach, a natural question is how sample weights are assigned across samples with varying cognitive levels. In Fig. 4, we collected all calculated sample weights and grouped them by corresponding cognitive score. Again, cognitive scores are not accessible during the training phase. On the ADReSS corpus, the weights roughly correlate with the MMSE scores, but not that well. We attribute this to the dispersion of sample groups in this dataset, where most groups contain fewer than 10 samples and can hardly guarantee statistical stability. Nevertheless, samples with MMSE scores below 15 tend to be assigned notably higher weights. In contrast, samples in the CUMV-R dataset display more distinct negative correlations – the less frequent groups are assigned larger weights. These findings imply that the InRe approach attempts to regulate the models’ training to focus more on AD instances within sparse groups. This behavior could explain the improvement in model sensitivity as presented in Table 1.
@@ 5 Conclusions
This work explores the issue of within-class variation in Alzheimer’s Disease (AD) detection. We posit that binary classification may overlook two crucial aspects: within-class heterogeneity (WCH) and instance-level imbalance (ILI). We further introduce two simple yet effective methods to address these problems: soft target distillation (SoTD) and instance-level re-balancing (InRe). Experiments on ADReSS and CU-MARVEL corpora demonstrated their advantages in detection performance. Future work will explore the combination of the SoTD and InRe methods, and investigate the incorporation of cognitive scores during model training.

### v3

# On the Within-class Variation Issue in Alzheimer's Disease Detection

@@ Abstract
Alzheimer's Disease (AD) detection employs machine learning classification models to distinguish between individuals with AD and those without. Different from conventional classification tasks, we identify within-class variation as a critical challenge in AD detection: individuals with AD exhibit a spectrum of cognitive impairments. Therefore, simplistic binary AD classification may overlook two crucial aspects: within-class heterogeneity and instance-level imbalance. In this work, we found using a sample score estimator can generate sample-specific soft scores aligning with cognitive scores. We subsequently propose two simple yet effective methods: Soft Target Distillation (SoTD) and Instance-level Re-balancing (InRe), targeting two problems respectively. Based on the ADReSS and CU-MARVEL corpora, we demonstrated and analyzed the advantages of the proposed approaches in detection performance. These findings provide insights for developing robust and reliable AD detection models.
@@
Alzheimer's disease, neurocognitive disorder, within-class variations, AD detection, dementia, healthcare1 Introduction
Neurocognitive disorders (NCD) such as Alzheimer's Disease (AD), present a substantial and growing challenge within the aging population, characterized by progressive cognitive decline across multiple domains, including memory, attention, and executive function [1]. For timely intervention and management, in-person clinical assessments have been the primary protocol for screening AD patients in populations, where participants are examined using specially designed assessment tasks to test potential abnormal declines in cognitive abilities[2, 3, 4]. In contrast to traditional on-set assessments, recent advancements in machine learning technologies have facilitated speech-based automatic AD detection as a promising screening approach, with the advantages of being scalable, accessible, and cost-efficient.
@@
Beyond the challenge of feature extraction and representation learning, we propose that AD detection faces unique difficulties inherent to the nature of Alzheimer's Disease. Medical literature [21, 22, 23] establishes AD as a degenerative disorder characterized by a continuum of pathophysiological changes, resulting in gradual cognitive and functional decline. This spectrum of cognitive performance among AD patients is illustrated in Fig. 1(a). Consequently, under standard classification modeling, samples with the same label may exhibit varying degrees of cognitive impairment patterns, which could lead to considerable variability in recognition features. This distinguishes AD detection from classification tasks such as image classification, which typically involve more consistent features under clearly defined categories. An ideal solution would be to model AD detection by regression or multi-way classification to capture this continuum. However, granular labels are rarely provided for most cognitive assessment tasks, especially for open-source datasets. While reference continuous scores are occasionally available (such as Mini-Mental State Examination scores), these scores are obtained from different tests that are distinct from the assessment in the data and can hardly be regarded as gold standard for supervision. Therefore, binary classification remains the prevailing method for Alzheimer's Disease (AD) detection.
This work takes a first step towards the within-class variation (WCV) issue in AD detection. We inspect the sample variability beyond binary labels using the English ADReSS[24] and Cantonese CU-MARVEL [6] datasets (see Fig. 1 (b) and Fig. 1 (c)). Even with comparable class members, AD samples functioning a larger variance in cognitive function compared to healthy samples. Given these findings, we propose that the conventional binary classification paradigm overlooks two critical aspects: a) within-class heterogeneity (WCH): samples with varying AD severity are assigned to the same class, potentially inhibiting the model's sensitivity to certain changes in cognitive function; and b) instance-level imbalance (ILI): the frequency of varying-severity samples is imbalanced even with balanced class size, thereby introducing potential bias. A key challenge in tackling these issues is the lack of informative instance-level labels indicating sample severity. Accordingly, this work explored sample score estimation that distills proxy soft labels from ensemble models with only hard label supervision. Building upon sample score estimation, we subsequently propose two approaches addressing WCH and ILI respectively: Soft Target Distillation (SoTD) and Instance-level Re-balancing (InRe). SoTD approach leverages the label refinery approach[25] to train classifiers using informative soft label supervision. And InRe approach re-weights imbalanced samples at instance level using log posterior ratios and inverse kernel density.
@@ 4.1 Validate estimated sample scores
An effective sample score estimation serves as a prerequisite for modeling within-class heterogeneity and imbalance using SoTD and InRe approaches. We first generate estimated scores for every sample in two datasets and observe their correlation with participants' cognitive scores (i.e., MMSE scores in ADReSS; MocA scores in CU-MARVEL). It is important to note that the cognitive scores referenced here should be viewed as assistant labels that indicate participants' cognitive ability in certain aspects, whereas the hard labels represent ground-truth labels for AD. As drawn in Fig. 3, we can observe that although the cognitive score labels were not seen during training, the estimated sample scores exhibit alignment with cognitive scores. Their Pearson Correlation is and in the two datasets respectively. Correspondingly, as used in the proposed SoTD method, guiding classifier training with these soft targets would take into account more comprehensive cognitive abilities of participants compared to relying on binary hard labels. Although we found some soft targets show ``noisy'' deviations from cognitive scores, we contend that the model overall benefits from these soft targets rather than just binary labels with values of 1 and 0.
@@ 4.2 AD detection performance
| InRe* | .5518 | .5881 | ||
Table 1 presents our experimental results on ADReSS and CUMV-R corpus. To reduce randomness, we performed 10-fold cross-validation for 20 random runs and reported the averaged results. On the ADReSS corpus, the SoTD approach outperformed the baseline and ensemble systems across all 4 metrics. This demonstrates general improvement in AD classification. The InRE approach also led to improvement in the balanced accuracy and F1 metrics, while remarkably enhancing the recall rates. This aligns with our expectation because the InRe approach could emphasize the infrequent positive (AD) samples in the dataset (as shown in Fig. 1 (b)), thereby improving the model's sensitivity. It is worth noting that this contribution differs from conventional class re-balancing methods since the number of positive and negative samples is equal in the training set. Thus, we contend that the advantage of the InRe approach stems its awareness of within-class imbalance. The resampling approach was not carried out in this dataset, as this dataset is already balanced and the results will be the same as the baseline.
The CUMV-R dataset poses a more challenging scenario for several reasons: (a) it contains more variations in the cognitive scores; (b) positive and negative samples are more imbalanced (ratio=1.51); and (c) Cantonese BERT models might not be as well-developed as English models. This is particularly evident in the low recall rates. The SoTD approach outperform both baseline and ensembling methods by balancing the model's predictions, evidenced by the improved recall rate. The resampling and InRe methods both show further improvement in recall rate, while the InRe method improved the models' sensitivity by a large margin () and therefore remarkably improved the overall F1 score ().
@@ 4.3 Inspect sample weights in InRe approach
To gain deeper insight into the InRe approach, a natural question is how sample weights are assigned across samples with varying cognitive levels. In Fig. 4, we collected all calculated sample weights and grouped them by corresponding cognitive score. Again, cognitive scores are not accessible during the training phase. On the ADReSS corpus, the weights roughly correlate with the MMSE scores, but not that well. We attribute this to the dispersion of sample groups in this dataset, where most groups contain fewer than 10 samples and can hardly guarantee statistical stability. Nevertheless, samples with MMSE scores below 15 tend to be assigned notably higher weights. In contrast, samples in the CUMV-R dataset display more distinct negative correlations – the less frequent groups are assigned larger weights. These findings imply that the InRe approach attempts to regulate the models' training to focus more on AD instances within sparse groups. This behavior could explain the improvement in model sensitivity as presented in Table 1.
@@ 4.4 Effect of bandwidths in InRe
4.5 Sample logits distribution in SoTD 333This section was omitted from the conference version due to page constraints.
To further prob the inner working of the SoTD approach, Fig. 5 depicts the logits (pre-softmax values) distributions of the baseline and proposed SoTD models, as a comparison of their discriminability of within-class sub-groups. Specifically, this experiment was conducted on the CUMV-R dataset, which provides additional 3-way labels: Healthy Control, Mild AD, and Major AD. In both systems, the models are trained on binary classification where we group "Mild AD" and "Major AD" as one "AD" group, in contrast to the "Healthy Control" group. Therefore, the "Mild AD" and "Major AD" are implicit within-class subgroups to the models. During inference, the logits of all three groups are recorded and their distribution are shown in Fig. 5. For the baseline model, we found that the logits of subgroups are almost fully overlapped, which showcases our concern about within-class variation: a naive AD detection model trained with binary labels can hardly be aware of the inherent within-class difference of AD patients. As a comparison, the subgroups can be relatively better distinguished by the SoTD model, as the purple and yellow distributions are less overlapped. Despite certain progress was been made, we still hope to emphasize that most of the samples in the two subgroups still overlapped, therefore within-class variation is still a challenging problem for the AD detection task.
@@ 5 Conclusions
This work explores the issue of within-class variation in Alzheimer's Disease (AD) detection. We posit that binary classification may overlook two crucial aspects: within-class heterogeneity (WCH) and instance-level imbalance (ILI). We further introduce two simple yet effective methods to address these problems: soft target distillation (SoTD) and instance-level re-balancing (InRe). Experiments on ADReSS and CU-MARVEL corpora demonstrated their advantages in detection performance. Future work will explore the combination of the SoTD and InRe methods, and investigate the incorporation of cognitive scores during model training.